Evaluate an investment projects and choose best investment option. Bank deposit rate 6% per annum:

1) Investment 2000, gives a profit at the end of the 1st year 200, 2nd year.b. 500, 3.g.b. EUR 1200;

2) a bon (4 coupons) with 500 EUR face value. The annual benefit (coupon payment) is 50 EUR. The term of the bond is 4 years.


1
Expert's answer
2021-01-13T10:23:46-0500

If bank deposit rate 6% per annum, then:

1) "NPV = -2,000 + 200\/1.06 + 500\/1.06^2 + 1,200\/1.06^3 = -358.78."

2) "NPV = -2,000 + 200\/1.06 + 200\/1.06^2 + 200\/1.06^3 + 200\/1.06^4 = -1,306.98."

So, the first option is better.